# CAMAR: Continuous Actions Multi-Agent Routing

### v1

@@ Introduction
Multi-agent reinforcement learning (MARL) has shown strong results in cooperative and competitive settings. Many recent studies explore how MARL can solve various tasks, including navigation and coordination, even in complex environments de2020independent ; bettini2023heterogeneous ; yu2022surprising ; damani2021primal ; skrynnik2024learn ; andreychuk2025mapf . However, current MARL benchmarks do not fully match the needs of real-world applications such as robotics. Among the many MARL scenarios, navigation and path planning are especially important for robotic systems. Yet, most existing benchmarks simplify navigation into grid worlds and discrete actions that fail to capture the smooth motion and collision avoidance that robots require.
@@
In addition, learnable methods have recently become the norm for pathfinding tasks because they can handle many agents at once angulo2022policy ; skrynnik2024learn ; andreychuk2025mapf ; damani2021primal . At the same time, non-learnable approaches struggle with computation as the number of agents grows. Therefore, this work focuses on learnable approaches for path finding in continuous multi-agent settings.
Robots move in continuous space and follow real physics, so they need to plan smooth paths that avoid walls and other robots simultaneously. The task of multi-robot navigation and avoiding collisions with dynamic obstacles is fundamental in robotics and has been studied for a long time chen2024survey ; wang2020mobile ; angulo2022policy ; lehoux2024multi . At the same time, most of the work uses its own specific environments and tasks, primarily built using slow simulators such as Gazebo [^1] https://gazebosim.org/home and Isaac Sim [^2] https://developer.nvidia.com/isaac/sim. To do this well, we need environments that combine continuous movement with multi-agent interaction. These environments must run fast, use realistic physics, and support many agents learning to navigate without collisions. High speed lets researchers run many experiments and test methods quickly. They should also include built-in benchmarks and evaluation tasks so researchers can compare different methods under the same conditions easily.
We have identified three main limitations in existing MARL environments. First, many use discrete environments, which cannot represent the smooth movements and physical dynamics important in real applications papoudakis2020benchmarking ; skrynnik2022pogema ; skrynnik2025pogema . Second, while some benchmarks do support continuous environments and complex tasks, they do not scale well when the number of agents and obstacles becomes large bettini2022vmas . Third, some environments can scale and use continuous states and actions, but their tasks are too simple and do not challenge agents to develop advanced policies ellis2023smacv2 ; lowe2017multi ; rutherford2023jaxmarl .
Because of these limitations, a new MARL benchmark for robotic navigation tasks that combines three key features is clearly needed: It should support continuous state and action spaces, handle hundreds or thousands of agents, and include tasks complex enough to reflect the challenges of real-world problems.
@@
To sum up, our benchmark builds on the strengths of existing environments and aims to complement ongoing efforts in MARL research. Over the years, a range of benchmarks has helped researchers tackle challenges in coordination, planning, and navigation. Well-known examples include SMAC samvelyan2019starcraft ; ellis2023smacv2 ; rutherford2023jaxmarl , Jumanji bonnet2023jumanji , POGEMA skrynnik2022pogema ; skrynnik2025pogema , MPE mordatch2017emergence ; lowe2017multi , and VMAS, each of which introduces useful features for different kinds of tasks.
SMAC samvelyan2019starcraft ; ellis2023smacv2 ; rutherford2023jaxmarl is popular for testing strategic decision-making, but it uses discrete actions and does not scale well for large environments. Jumanji bonnet2023jumanji supports GPU acceleration and procedural generation, but its focus is not on navigation or planning. POGEMA skrynnik2022pogema ; skrynnik2025pogema is strong in grid-based navigation and procedural tasks with many agents, but it does not use continuous states or actions, which are important in robotics.
MPE mordatch2017emergence ; lowe2017multi has played an important role in early MARL research, but it cannot scale to hundreds of agents efficiently. VMAS bettini2022vmas builds on MPE mordatch2017emergence ; lowe2017multi by adding its physics and continuous dynamics, making it better for robotics. However, it can still be slow and difficult to scale to larger agent populations or complex maps.
Finally, popular robotics simulators such as Gazebo koenig2004design , Webots michel2004cyberbotics and ARGoS pinciroli2012argos offer continuous observations and actions with realistic physics, but they do not provide the ability for fast training in multi-agent scenarios. This gap shows the need for a new simulation environment that brings together high performance and support for large-scale multi-agent navigation and planning. For more detailed analysis of each environment see Appendix H.
@@
Waterworld (SISL) gupta2017cooperative link
RWare papoudakis2020benchmarking link
@@
RWare (Pufferlib) suarez2024pufferlib link
Trash Pickup (Pufferlib) suarez2024pufferlib link
@@
JaxNav (JaxMARL) rutherford2024no link
Nocturne vinitsky2022nocturne link
@@
Continuous Observations and Actions
Robots usually operate in continuous state and action spaces, so it’s important for benchmarks to reflect these conditions. Environments like VMAS and MPE mordatch2017emergence ; lowe2017multi use continuous actions and observations, while others, like Trash Pickup suarez2024pufferlib or POGEMA skrynnik2025pogema , use discrete representations, which limit realism in robotics simulations.
GPU Support
In multi-agent environments, each agent often has its observations and rewards. This can lead to heavy data transfers between the CPU and the GPU, especially when training deep RL models. Benchmarks like Jumanji bonnet2023jumanji and SMAX rutherford2023jaxmarl are initially built with GPU acceleration, helping to reduce training time. In contrast, many older environments like SMAC samvelyan2019starcraft ; ellis2023smacv2 and MPE mordatch2017emergence ; lowe2017multi do not support GPU-based simulation.
Scalability >500 Agents
When the number of agents grows, decision-making becomes harder. A good benchmark should scale well to hundreds or thousands of agents. Environments like POGEMA skrynnik2022pogema ; skrynnik2025pogema and Trash Pickup suarez2024pufferlib can run with thousands and even millions of agents. Others, like VMAS, MPE mordatch2017emergence ; lowe2017multi , and SMAC samvelyan2019starcraft ; ellis2023smacv2 , are limited to much smaller groups.
Partially Observable
In most real-world scenarios, agents see only part of the environment. This is known as partial observability and is a common feature in almost every RL environment. SMAC samvelyan2019starcraft ; ellis2023smacv2 , RWare papoudakis2020benchmarking , and POGEMA skrynnik2022pogema ; skrynnik2025pogema support this feature by limiting each agent’s view.
Heterogeneous Agents
In real life, agents and robots can have different sensors, shapes, or goals. Benchmarks like VMAS and SMAC samvelyan2019starcraft ; ellis2023smacv2 allow agents to behave differently, making cooperation more complex. Others, like POGEMA skrynnik2022pogema ; skrynnik2025pogema or RWare papoudakis2020benchmarking , usually involve homogeneous agents.
Performance >10K Steps/s
High simulation speed is essential when agents need millions of interactions to learn. For example, the Pufferlib suarez2024pufferlib environment can run with a speed of up to 1M steps per second. This allows researchers to train and develop models more quickly. Environments like SMAC samvelyan2019starcraft ; ellis2023smacv2 or VMAS can be slower, especially when running with many agents.
Python Based
Using Python makes it easier for researchers to understand and modify the environment. SMAX rutherford2023jaxmarl , VMAS, and POGEMA skrynnik2022pogema ; skrynnik2025pogema are implemented entirely in Python. This helps with faster development and integration into learning frameworks.
Procedural Generation
Procedural generation helps create diverse tasks that reduce the chance of overfitting. POGEMA skrynnik2022pogema ; skrynnik2025pogema and Jumanji bonnet2023jumanji use this method to generate complex tasks automatically. Other environments, like RWare papoudakis2020benchmarking , use fixed layouts that limit diversity.
Requires Generalization
Only a few environments offer separate training and testing tasks, making it difficult to test whether agents generalize well to new situations. POGEMA skrynnik2022pogema ; skrynnik2025pogema and SMACv2 ellis2023smacv2 include test scenarios that allow researchers to check generalization. Many others use the same tasks during both training and testing.
3 SPS decreases gracefully with many agents and obstacles.
Evaluation Protocols
To compare algorithms fairly, we need well-defined test cases and metrics. Benchmarks like SMACv2 ellis2023smacv2 and Jumanji bonnet2023jumanji include evaluation protocols to make results reproducible and meaningful. Many other environments do not have standard evaluation tools or test setups.
Tests & CI
PyPi Listed
When an environment is available on PyPi, it becomes easier for others to install and use. Benchmarks like Jumanji bonnet2023jumanji , and Trash Pickup suarez2024pufferlib are available on PyPi. This lowers the barrier to entry and helps increase adoption in the research community.
@@
$\begin{cases}u_{i}(t)=\text{clip}(u_{i}^{a}(t),-\texttt{max\_u},\texttt{max\_u})\\ \omega_{i}(t)=\text{clip}(\omega_{i}^{a}(t),-\texttt{max\_w},\texttt{max\_w})\\ \vec{v}_{i}(t)=[u_{i}(t)\cos(\theta_{i}(t));u_{i}(t)\sin(\theta_{i}(t)]\\ \vec{\text{pos}}_{i}(t+dt)=\vec{\text{pos}}_{i}(t)+\vec{v}_{i}(t)dt\\ \theta_{i}(t+dt)=\theta_{i}(t)+\omega_{i}(t)dt\end{cases}$
@@
(a) random_grid
(b) labmaze_grid
(c) movingai
(d) caves_cont
(e) string_grid
(f) batched_string_grid
@@
In multi-agent pathfinding research, it is common to use metrics such as success rate (SR), flowtime (FT), makespan (MS), and coordination (CO) to compare different methods. We adopt these metrics in our work and define them more formally as follows in the equation below:
$\begin{cases}\mathrm{SR}=\frac{1}{N}\sum\limits_{i=1}^{N}\mathbf{1}\{\|\vec{x}_{i}(T)-\vec{x}_{\text{goal}}\|\leq R_{\text{g}}\};\\ \mathrm{FT}=\frac{1}{N}\sum\limits_{i=1}^{N}t_{i};\\ \mathrm{MS}=\max\limits_{i=1,\dots,N}t_{i};\\ \mathrm{CO}=1-\frac{C}{N\times T}\end{cases}$
Here, $N$ is the number of agents, $T$ is the maximum episode length, $t_{i}$ is the time step when agent $i$ first reaches its goal, $C$ is the total number of collisions over the episode.
@@
We evaluate six MARL algorithms: IPPO, MAPPO, IDDPG, MADDPG, ISAC haarnoja2018soft , and MASAC. In addition, we include two non-learnable baselines, RRT+PD and RRT*+PD, and two hybrid methods, RRT*+IPPO and RRT*+MAPPO.
All methods are evaluated on two procedurally generated map types: random_grid and labmaze_grid, each with 6 versions varying in obstacle density and agent count (8 or 32). For labmaze_grid, an additional connection probability from 0.4 to 1.0 tests different maze complexities. Generation details for training and evaluation maps are available in the Appendix D.
@@
The main results are presented in Table 2, which reports success rate (SR), flowtime (FT), makespan (MS), and coordination (CO) for each algorithm on two types of maps.
On the random_grid map, MAPPO achieves the highest SR, with good FT and full CO. Its centralized critic helps agents plan jointly and share useful information. RRT*+MAPPO shows similar SR but performs better in FT and MS, which shows that adding planning support improves movement efficiency.
RRT*+IPPO also improves over IPPO alone, although the gains are smaller. RRT*+PD achieves high SR without learning, but it performs worse in CO. This is expected because it plans for each agent independently. The planner uses full map knowledge but does not consider other agents during execution. As a result, agents sometimes collide in narrow areas.
RRT* with Off-Policy Algorithms
We also evaluate hybrid methods that combine RRT* with off-policy MARL algorithms. As shown in Table 2, the results are mixed. Some methods benefit from additional planning information. For example, RRT*+ISAC slightly improves over ISAC alone in both success rate (SR) and coordination (CO), especially in the random_grid setting. RRT*+IDDPG also shows moderate improvement, achieving better SR and lower flowtime (FT) than IDDPG.
However, in other cases, the integration leads to weaker results. RRT*+MASAC and RRT*+MADDPG perform poorly, with SR staying below 0.05 in both maps. One possible reason is the increased input size to the models. Off-policy algorithms with centralized critics, such as MASAC and MADDPG, must process long input vectors that include observations and RRT* features from all agents. This can make training unstable and reduce generalization.
Overall, while combining RRT* with off-policy methods can help in some settings, the results suggest that better integration strategies or more expressive policy architectures may be needed to fully benefit from hybrid approaches.
Among MARL methods, IPPO and IDDPG perform well in SR, but their FT is worse than MAPPO. ISAC shows high CO but low SR, possibly due to its focus on smooth motion and risk reduction. MASAC and MADDPG perform poorly. In particular, MADDPG shows very low SR and fails to explore even in the simpler tasks. We investigate these issues in more detail in the Appendix D.
The labmaze_grid map is more challenging. Narrow passages make reward signals sparse. In this setting, SR drops across all MARL methods. RRT*+PD performs the best in SR, showing the value of full-path planning when learning signals are limited. However, its CO score remains lower than MARL baselines.
The simpler RRT+PD baseline performs worse than RRT*+PD across all metrics. It produces longer and less efficient paths because RRT does not optimize for path cost and often generates jagged routes. On the labmaze_grid, RRT+PD still performs better in SR than many learning-based methods, but with higher FT and lower CO. This result shows that even basic planning methods can outperform MARL baselines in sparse reward environments, but their inability to consider other agents during execution limits coordination and overall efficiency.
These findings highlight a trade-off. Planning methods like RRT*+PD solve complex tasks more reliably, while learning-based policies handle coordination better. To explore this further, we evaluated hybrid methods. These approaches combine RRT* samples with MARL observations to guide exploration. However, results show only small gains or similar performance. One possible reason is the size of the observation vector, which may be too large for the simple policy network to process effectively. Future work may improve integration with better architectures. See Appendix D for more detailed analysis of performance of algorithms.
@@
To show CAMAR’s support for heterogeneous agents, we modify a simple coordination task called give_way. In this new version, two agents must pass through a narrow corridor, but only the smaller red agent can fit into the central chamber. The larger blue agent cannot enter this area and must wait for the other to pass.
We evaluate both standard and heterogeneous versions of IPPO and MAPPO. In the heterogeneous versions (HetIPPO and HetMAPPO), agents use separate policy models. Results in Fig. 6(b) show that HetIPPO performs better than its shared-policy version. However, HetMAPPO fails, likely because the centralized critic struggles with large and diverse input spaces.
@@
This experiment shows that CAMAR supports agents with different sizes and models. It can be used to study coordination and policy learning in heterogeneous teams. More details and discussion are provided in the Appendix D.
@@
(c) CAMAR
(d) VMAS
We evaluate the scalability of our environment by measuring how the simulation speed changes when we increase the number of parallel environments, agents, and obstacles. Fig. 7 and  8 show the results of these experiments.
Fig. 7 shows a snapshot of the CAMAR environment with 32 agents on the $20\times 20$ random_grid map. Agents move smoothly while avoiding the 120 fixed obstacles. Fig. 7 shows the same scenario in VMAS bettini2022vmas .
@@
First, we measure the SPS as we increase the number of parallel environments from $5$ to $7000$, keeping $32$ agents fixed. VMAS bettini2022vmas scales almost linearly but reaches less than 10 00010\,000 SPS at $7000$ environments. In contrast, CAMAR scales quickly up to $2000$ environments and then maintains around 50 00050\,000 SPS even as more are added. This shows that CAMAR can simulate many parallel tasks without losing speed, which is important for fast training.
@@
Next, we fix the number of environments at $2000$ and vary the agent count from $4$ to $128$. CAMAR keeps over 100 000100\,000 SPS with fewer than $16$ agents and remains above 10 00010\,000 SPS even with $128$ agents. VMAS bettini2022vmas starts near 20 00020\,000 SPS with $4$ agents but drops sharply to $500$ SPS at $128$ agents. In this comparison, CAMAR is up to 2020x faster than VMAS when many agents act at once.
Overall, CAMAR delivers much higher throughput and keeps strong performance as the number of agents increases. VMAS can boost its speed by adding more parallel environments, but it would need about 50 00050\,000 parallel environments to match the throughput that CAMAR provides at $2000$ environments. We also compare CAMAR scalability with other modern MARL environments, which can be found in the Appendix E.
Figure 8: Scalability of CAMAR with increasing numbers of agents and obstacles under more extreme conditions. Despite a decrease in Steps per Second (SPS) as the number of agents grows to $800$, CAMAR maintains approximately $1400$ SPS, demonstrating its capability to simulate large multi-agent teams. Additionally, although SPS decreases, the total number of gathered observations remains high, as it scales proportionally with the number of agents.
We also study CAMAR under more extreme conditions. In Fig. 8, we increase the number of agents up to $800$ while keeping the number of circle obstacles fixed at $4160$. As the number of agents grows, the simulation speed drops, but our environment still maintains about $1400$ SPS with $800$ agents. This result shows that CAMAR can simulate very large multi-agent teams and still produce enough experience for learning. It also means that the total number of observations remains high, because each agent generates one observation per step.
@@
Finally, we test how the number of obstacles affects scalability. We fix the number of agents at $32$ and increase the number of obstacles up to $9920$. The results, shown in Fig. 8, demonstrate that although SPS decreases with more obstacles, CAMAR still achieves about 15 40015\,400 SPS at the highest tested complexity. This indicates that CAMAR remains robust even in very cluttered environments.

### v2

@@ Introduction
Multi-agent reinforcement learning (MARL) has shown strong results in cooperative and competitive settings, and many recent studies explore how MARL can solve tasks that need coordination in complex environments de2020independent ; bettini2023heterogeneous ; yu2022surprising ; damani2021primal ; skrynnik2024learn ; andreychuk2025mapf . One important group of such tasks is multi-agent pathfinding (MAPF), where several agents must reach their goals without collisions. Classic MAPF is usually studied on discrete grids, but real robots move in continuous space, follow dynamics, and must plan smooth paths. This continuous version of MAPF is important for many domains such as warehouse logistics, drone swarm coordination, and other systems where many robots must move safely and efficiently.
@@
Learnable methods have recently become effective for MAPF and cooperative navigation angulo2022policy ; skrynnik2024learn ; andreychuk2025mapf ; damani2021primal . However, many existing environments still use grid worlds or discrete actions that do not match real robot behavior. In practice, robots must plan and coordinate while avoiding both static and dynamic obstacles, a long-standing problem in robotics chen2024survey ; wang2020mobile ; angulo2022policy ; lehoux2024multi . Many MARL environments simplify this problem too much by using low-dimensional maps or unrealistic movement models.
High-fidelity simulators like Gazebo [^1] https://gazebosim.org/home, Isaac Sim [^2] https://developer.nvidia.com/isaac/sim, AirSim shah2017airsim , and Flightmare song2021flightmare offer realistic physics, but they focus on robot control and perception, not on large-scale MARL. These tools often run slower and cannot simulate hundreds of agents at once. To study large-scale continuous navigation, we need environments that combine high speed with continuous dynamics, run efficiently on GPU, and allow testing both learned and planning-based methods in a single benchmark.
We identify three main gaps in current MARL environments. First, many use discrete action spaces that cannot represent smooth motion papoudakis2020benchmarking ; skrynnik2022pogema ; skrynnik2025pogema . Second, while some environments do support continuous states and actions, they do not scale to large numbers of agents and obstacles bettini2022vmas . Third, other environments can scale but offer simple tasks that do not require strong coordination or realistic navigation skills ellis2023smacv2 ; lowe2017multi ; rutherford2023jaxmarl .
@@
To compare existing MARL environments, we evaluate them in Table 1 across key features such as continuous control, GPU support, scalability, and usability. An extended version of the Table 1 with descriptions of comparison features and detailed analysis of each environment are provided in the Appendix H.
Prominent benchmarks include SMAC samvelyan2019starcraft ; ellis2023smacv2 ; rutherford2023jaxmarl , Jumanji bonnet2023jumanji , POGEMA skrynnik2022pogema ; skrynnik2025pogema , MPE mordatch2017emergence ; lowe2017multi , and VMAS bettini2022vmas . SMAC enables testing of strategic behavior but uses discrete actions and scales poorly. Jumanji supports GPUs and procedural generation but is not focused on navigation. POGEMA handles large-agent navigation with procedural maps but lacks continuous control. MPE, while foundational, does not scale to many agents; VMAS adds physical realism but still struggles with performance and scalability.
Many environments lack evaluation protocols and suffer from slow training due to CPU-GPU bottlenecks. Simulators like Gazebo koenig2004design , Webots michel2004cyberbotics , and ARGoS pinciroli2012argos offer realistic continuous dynamics but are not optimized for efficient MARL training. These gaps highlight the need for a new benchmark supporting scalable, high-performance multi-agent learning.
@@
3 SPS decreases gracefully with many agents and obstacles.
$\begin{cases}u_{i}(t)=\text{clip}(u_{i}^{a}(t),-\texttt{max\_u},\texttt{max\_u})\\ \omega_{i}(t)=\text{clip}(\omega_{i}^{a}(t),-\texttt{max\_w},\texttt{max\_w})\\ \vec{v}_{i}(t)=[u_{i}(t)\cos(\theta_{i}(t));u_{i}(t)\sin(\theta_{i}(t)]+\frac{\vec{f}^{c}_{i}(t)}{m}dt\\ \vec{\text{pos}}_{i}(t+dt)=\vec{\text{pos}}_{i}(t)+\vec{v}_{i}(t)dt\\ \theta_{i}(t+dt)=\theta_{i}(t)+\omega_{i}(t)dt\end{cases}$
@@
(a) random_grid
(b) labmaze_grid
(c) movingai
(d) caves_cont
(e) string_grid
(f) batched_string_grid
@@
The evaluation metrics are defined as follows: the success rate is $\mathrm{SR}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{\|\vec{x}_{i}(T)-\vec{x}_{\text{goal}}\|\leq R_{\text{g}}\}$; the flowtime is $\mathrm{FT}=\frac{1}{N}\sum_{i=1}^{N}t_{i}$; the makespan is $\mathrm{MS}=\max_{i=1,\dots,N}t_{i}$; and the coordination is $\mathrm{CO}=1-\frac{C}{N\times T}$. Here, $N$ is the number of agents, $T$ is the maximum episode length, $t_{i}$ is the time step when agent $i$ first reaches its goal, $t_{i}=T$ for unfinished agents, $C$ is the total per-agent-per-timestep number of collisions over the episode.
@@
We evaluate 6 MARL algorithms: IPPO, MAPPO, IDDPG, MADDPG, ISAC haarnoja2018soft , and MASAC. In addition, we include 2 non-learnable baselines, RRT+PD and RRT*+PD, and 6 hybrid methods: RRT*+IPPO, RRT*+MAPPO, RRT*+IDDPG, RRT+MADDPG, RRT*+ISAC, and RRT*+MASAC.
All methods are evaluated on two procedurally generated map types: random_grid and labmaze_grid, each with 6 versions that vary in obstacle density and agent count (8 or 32) [^6] Our current Medium-tier protocol includes tasks with 8 and 32 agents, but we plan to extend it in future versions to support larger agent populations and more complex settings as methods advance.. For labmaze_grid, an additional connection probability ranging from 0.4 to 1.0 is used to test different maze complexities. Generation details for all training and evaluation maps are provided in Appendix D.
@@
The main results are shown in Table 2. On the random_grid map, MAPPO reaches the highest success rate with strong coordination. Adding planning improves efficiency. RRT*+MAPPO gives faster routes than MAPPO, and RRT*+IPPO improves over IPPO. The classic RRT*+PD baseline reaches high success without learning, but shows low coordination because it plans for each agent alone.
For off-policy algorithms, results are mixed. RRT* improves ISAC and IDDPG slightly, but weakens MASAC and MADDPG. These methods use a centralized critic and must process long input vectors that include RRT* features. This makes training unstable and harms generalization. Among MARL baselines, IPPO and IDDPG reach good success but slower flowtime than MAPPO. MASAC and MADDPG fail in both maps.
The labmaze_grid setting is harder due to narrow corridors and sparse rewards. All MARL baselines drop in success. RRT*+PD performs best, which shows the value of full-path planning when learning signals are weak. The simpler RRT+PD planner gives worse paths but still beats many MARL baselines in success.
More detailed analysis of these results and hybrid methods is given in Appendix D.
@@
To demonstrate support for heterogeneous teams, we extend the give_way task so that two agents must pass through a narrow corridor where only the smaller red agent can enter the central chamber. The larger blue agent must wait.
We compare IPPO and MAPPO with shared policies to their heterogeneous versions, where each agent has its own model. As shown in Fig. 6(b), HetIPPO performs better, but HetMAPPO fails, likely because the centralized critic cannot handle the larger and more diverse input space.
@@
This experiment shows that CAMAR can model agents with different sizes and abilities and is suitable for studying coordination in heterogeneous teams. More details are in Appendix D.
@@
We study how CAMAR scales when we increase the number of parallel environments, agents, and obstacles. The full results are shown in Fig. 7. Visual examples of CAMAR and VMAS running the same scenario are provided in the appendix E.
@@
We first fix the number of agents at 32 and measure the simulation speed when we increase the number of parallel environments from $5$ to 60006000+. VMAS scales roughly linearly but stays below 10 00010\,000 steps per second (SPS) at $6000$ environments. In contrast, CAMAR rises quickly to about $1000$ environments and then stays close to 50 00050\,000 SPS even as we add more parallel environments. This shows that CAMAR can support fast and stable training with many vectorized environments.
@@
Next, we fix the number of environments at $2000$ and increase the number of agents from $4$ to $128$. CAMAR keeps more than 100 000100\,000 SPS when the agent count is below $16$ and remains above 10 00010\,000 SPS even at $128$ agents. VMAS begins near 20 00020\,000 SPS with $4$ agents but drops to about $500$ SPS at $128$ agents. In this setting, CAMAR is up to $20$ times faster when many agents act together.
(a) CAMAR vs VMAS with increasing number of environments, agents, and obstacles.
(b) CAMAR handles up to 800 agents, maintaining 1400 SPS and high observation throughput.
Figure 7: CAMAR achieves robust scalability in multi-agent simulation, surpassing VMAS in performance across varied settings and supporting efficient simulation of large agent populations with high observation rates.
We also test extreme conditions by keeping $4160$ fixed and increasing the number of agents up to $800$. The speed goes down as more agents are added, but CAMAR still maintains about $1400$ SPS at $800$ agents. Since every agent produces one observation per step, the total amount of data remains high.
@@
Finally, we test how obstacle count alone affects speed. We fix the number of agents at $32$ and increase the number of obstacles up to $9920$. CAMAR still reaches about 15 00015\,000 SPS in this most cluttered setting. This shows that the system remains robust even in very dense maps.
In summary, CAMAR achieves around 50 00050\,000 SPS on complex scenarios with many obstacles and $32$ agents. The speed depends mainly on the number of objects in the scene. In other tasks it reaches up to 161 000161\,000 SPS with $32$ agents and $960$ obstacles, and up to 562 000562\,000 SPS when simulating a single agent with $4160$ obstacles. These results confirm that CAMAR can operate above 100 000100\,000 SPS. Visualisations of the test scenes are shown in the appendix E. Further comparisons with other MARL environments are also given in the appendix E.
Acknowledgments
The study was supported by the Ministry of Economic Development of the Russian Federation (agreement No. 139-15-2025-013, dated June 20, 2025, IGK  [^00] C313925P4B0002).